OpenStudy (danise):

ALGEBRA!!!! 6,13,20,27,34 Write an explicit formula? an= Write a recursive formula? an= a1=

myininaya (myininaya):

you might notice there is a common difference

myininaya (myininaya):

for example what are the following equal to: 13-6=? 20-13=? 27-20=? 34-27=?

OpenStudy (danise):

7

myininaya (myininaya):

right and we know arithmetic sequences have this form: \[a_n=a_1+d(n-1)\] this an explicit form

myininaya (myininaya):

for a recursive form notice term-previous term=common difference